# Agents4PLC: Automating Closed-loop PLC Code Generation and Verification in Industrial Control Systems using LLM-based Agents

### v1

@@ IV Experimental Evaluation
Benchmark Construction. To accurately evaluate our Agents4PLC system, we constructed the first benchmark dataset focused on the task of generating Structured Text (ST) code from natural language specification, and we assess the correctness of code samples automatically through formal verification methods. This dataset comprises 23 programming tasks along with corresponding formal verification specifications, including 58 properties in easy set with 53 non-trivial property over 16 easy programming tasks and 43 in medium set with 38 non-trivial property over 7 medium programming tasks, where trivial property means ”assertion” property without corresponding assertion sentences in ST code for reference. These programming problems cover various aspects of industrial programming including Logical Control, Mathematical Operations, Real-time Monitoring, Process Control and other fields, which effectively simulate the genuine requirements found in industrial control systems.
@@
Evaluation Metrics. Following previous work[14], we employ the 1) pass rate, 2) syntax compilation success rate and 3) verification success rate as metrics to evaluate the effectiveness of code generation. The syntax compilation success rate serves as a preliminary validation of the syntactic correctness of the generated ST code. A high syntax compilation success rate indicates fewer syntax errors, thereby reducing subsequent debugging efforts. The pass rate is derived from the pass@k metric, where the model is considered successful if at least one of the k generated results can compile. A high pass rate reflects the effectiveness of the approach in generating executable ST code. We calculate it as the proportion of code segments that successfully compile out of the total number of generated code segments. The verifiable rate quantifies the proportion of the generated ST code segments that not only compile successfully but also adhere to the specified functional requirements of the PLC program, which is the most challenging task.
@@ V-A RQ1: Effectiveness study
In this experiment, we systematically compare our framework with the LLM4PLC framework based on several base LLMs for generating reliable ST code. Our evaluation is based on a set of benchmark cases designed to reflect varying levels of complexity, from relatively straightforward control sequences (labeled as “Easy”) to more sophisticated logic processes (labeled as “Medium”). Note that LLM4PLC is designed to be a half-automated framework with human interactions, we write an extra automation program to drive the components of the LLM4PLC framework. More experiment details are included in our site. [^2] https://hotbento.github.io/Agent4PLC/
@@
The experiment results is shown in Table I, which illustrates the performance of different frameworks based on different models across both “Easy” and “Medium” benchmark levels. The table presents the pass rates for both compilation and verification stages, with the format X Y Z%, where X represents the number of successful passes on corresponding metrics, Y indicates the total number of programming problems, and Z% denotes the pass rate percentage. For instance, LLM4PLC/GPT-4o achieves a syntax compilation pass rate of 14 16 87.5%, meaning the generated result of LLM4PLC on the GPT-4o model successfully compiles for 14 out of 16 programming problems, yielding an 87.5% success rate. We record syntax compilation success rate, verifiable rate and pass rate for both systems, where “verifiable rate” means the framework can generate verifiable model for at least 80% of the given properties, and “pass rate” means that least 80% of the generated code can compile.
@@
TABLE I: Compilation and Verification Metrics for Different Models
Benchmark Level Easy Problems Medium Problems
DeepSeek V2.5 13  16  81.3% 2  16  12.5% 2  16  12.5% 7  7  100.0% 0  7  0.0% 0  7  0.0%
GPT-4o 14  16  87.5% 0  16  0.0% 2  16  12.5% 4  7  57.1% 0  7  0.0% 0  7  0.0%
GPT-4o-mini 15  16  93.8% 0  16  0.0% 0  16  0.0% 4  7  57.1% 0  7  0.0% 0  7  0.0%
Agents4PLC CodeLlama 34B 5  16    31.3% 1  16  6.3% 2  16  12.5% 1  7  14.3% 0  7  0.0% 0  7  0.0%
DeepSeek V2.5 16  16  100.0% 7  16  43.8% 10  16  62.5% 7  7  100.0% 1  7  14.3% 3  7  42.9%
GPT-4o 16  16  100.0% 8  16  50.0% 11  16  68.8% 7  7  100.0% 2  7  28.6% 3  7  42.9%
GPT-4o-mini 16  16  100.0% 7  16  43.8% 8  16  50.0% 7  7  100.0% 0  7  0.0% 1  7  14.3%
ChatDev GPT-4o 7  16  43.8% 7  16  43.8% 7  16  43.8% 2  7  28.6% 1  7  14.3% 2  7  28.6%
TABLE II: Efficiency evaluation: ST code generation times over all attempts passing syntactic compilation
@@
TABLE III: Ablation Experiment with Designs on Coding and Fixing Agents
@@ V-B RQ2: Efficiency study
To evaluate the efficiency of our framework in generating validated test cases, we measured the time taken for each model to successfully generate a syntactically correct and semantically verified ST code. Considering in our validation process, PLCverif is not supported for efficiency evaluation, we categorize the results based on how many attempts it takes for each model to pass the syntactic compilation stage. The categories are as follows:
@@
The experimental results are presented in TableII, where the data format represents the number of generation attempts required / total cases passing syntactic compilation / ratio. The experimental setup is identical to that of Experiment 1.
@@ V-C RQ3: Ablation study
    Table III summarizes the findings from this ablation study, showcasing the ST code verification pass rates and compilation pass rates across various configurations. The configuration of experiment and table content is similar to those in effectiveness study, including metrics and data format. The configuration that on Easy Problem set, the setting combining One-shot and Syntax Hint yields the highest overall pass rates (68.8%) and verifiable rates (75.0%). However, the results for medium problems mirror the full configuration, One-shot + RAG + Syntax Hint configuration, demonstrate superior performance with 28.6% pass rate and 42.9% verifiable rate). The impact for each configuration analysis is:

### v2

@@ IV Experimental Evaluation
Benchmark Construction. To accurately evaluate our Agents4PLC system, we constructed the first benchmark dataset focused on the task of generating ST code from natural language specification, and we assess the correctness of code samples automatically through formal verification methods. This dataset comprises 23 programming tasks along with corresponding formal verification specifications, including 58 properties in easy set with 53 non-trivial property over 16 easy programming tasks and 43 in medium set with 38 non-trivial property over 7 medium programming tasks, where trivial property means ”assertion” property without corresponding assertion sentences in ST code for reference. These programming problems cover various aspects of industrial programming including Logical Control, Mathematical Operations, Real-time Monitoring, Process Control and other fields, which effectively simulate the genuine requirements found in industrial control systems.
@@
Evaluation Metrics. Following previous work[14], we employ the 1) pass rate, 2) syntax compilation success rate and 3) verification success rate (or verifiable rate) as metrics to evaluate the effectiveness of code generation. The syntax compilation success rate serves as a preliminary validation of the syntactic correctness of the generated ST code. A high syntax compilation success rate indicates fewer syntax errors, thereby reducing subsequent debugging efforts. The pass rate is derived from the pass@k metric, where the model is considered successful if at least one of the k generated results that not only compile successfully but also adhere to the specified functional requirements of the PLC program, which is the most challenging task. The verifiable rate quantifies the proportion of the generated ST code segments that can pass syntax check of target verification language, e.g. nuXmv or cbmc. A high verifiable rate reflects the effectiveness of the approach in generating executable code for verification. We calculate it as the proportion of code segments that successfully compile out of the total number of generated code segments.
@@ V-A RQ1: Effectiveness study
In this experiment, we systematically compare our framework with other code generation frameworks based on several base LLMs for generating reliable ST code. Our evaluation is based on a set of benchmark cases designed to reflect varying levels of complexity, from relatively straightforward control sequences (labeled as “Easy”) to more sophisticated logic processes (labeled as “Medium”). Note that LLM4PLC is designed to be a half-automated framework with human interactions, we write an extra automation program to drive the components of the LLM4PLC framework. More experiment details are included in our github link  [^2] https://github.com/Luoji-zju/Agents4PLC_release and our site. [^3] https://hotbento.github.io/Agent4PLC/
@@
The experiment results is shown in Table III, which illustrates the performance of different frameworks based on different models across both “Easy” and “Medium” benchmark levels. The table presents the pass rates for both compilation and verification stages, with the format X Y Z%, where X represents the number of successful passes on corresponding metrics, Y indicates the total number of programming problems, and Z% denotes the pass rate percentage. For instance, LLM4PLC/GPT-4o achieves a syntax compilation pass rate of 14 16 87.5%, meaning the generated result of LLM4PLC on the GPT-4o model successfully compiles for 14 out of 16 programming problems, yielding an 87.5% success rate. We record syntax compilation success rate, verifiable rate and pass rate for both systems, where “verifiable rate” means the framework can generate verifiable model for at least 80% of the given properties, and “pass rate” means that least 80% of the generated code can compile.
@@
TABLE I: Efficiency evaluation: ST code generation times over all attempts passing syntactic compilation
@@
TABLE II: Ablation Experiment with Designs on Coding and Fixing Agents
@@ V-B RQ2: Efficiency study
To evaluate the efficiency of our framework in generating verifiable st code, we measured the time taken for each model to successfully generate a syntactically correct and semantically verified ST code. Considering in our validation process, PLCverif is not supported for efficiency evaluation, we categorize the results based on how many attempts it takes for each model to pass the syntactic compilation stage. The categories are as follows:
@@
The experimental results are presented in Table I, where the data format represents the number of generation attempts required / total cases passing syntactic compilation / ratio. The experimental setup is identical to that of Experiment 1.
@@ V-C RQ3: Ablation study
    Table II summarizes the findings from this ablation study, showcasing the ST code verification pass rates and compilation pass rates across various configurations. The configuration of experiment and table content is similar to those in effectiveness study, including metrics and data format. The configuration that on Easy Problem set, the setting combining One-shot and Syntax Hint yields the highest overall pass rates (68.8%) and verifiable rates (75.0%). However, the results for medium problems mirror the full configuration, One-shot + RAG + Syntax Hint configuration, demonstrate superior performance with 28.6% pass rate and 42.9% verifiable rate). The impact for each configuration analysis is:
@@ V-D RQ4: Case study in practical control environment
TABLE III: Compilation and Verification Metrics for Different Models
Benchmark Level Easy Problems Medium Problems
Framework Base Model Syntax Compilation Verifiable Rate Pass Rate Syntax Compilation Verifiable Rate Pass Rate
DeepSeek V2.5 13  16  81.3% 2  16  12.5% 2  16  12.5% 7  7  100.0% 0  7  0.0% 0  7  0.0%
GPT-4o 14  16  87.5% 0  16  0.0% 2  16  12.5% 4  7  57.1% 0  7  0.0% 0  7  0.0%
GPT-4o-mini 15  16  93.8% 0  16  0.0% 0  16  0.0% 4  7  57.1% 0  7  0.0% 0  7  0.0%
Agents4PLC CodeLlama 34B 5  16    31.3% 2  16  12.5% 1  16  6.3% 1  7  14.3% 0  7  0.0% 0  7  0.0%
DeepSeek V2.5 16  16  100.0% 10  16  62.5% 7  16  43.8% 7  7  100.0% 3  7  42.9% 1  7  14.3%
GPT-4o 16  16  100.0% 11  16  68.8% 8  16  50.0% 7  7  100.0% 3  7  42.9% 2  7  28.6%
GPT-4o-mini 16  16  100.0% 8  16  50.0% 7  16  43.8% 7  7  100.0% 1  7  14.3% 0  7  0.0%
ChatDev GPT-4o 7  16  43.8% NA NA 2  7  28.6% NA NA